OpenStudy (anonymous):

he area of a circular sun spot is growing at a rate of 1,800 km2/s. How fast is the radius growing at the instant when it equals 5,000 km?

OpenStudy (anonymous):

\[A=\pi r^2\] Differentiating with respect to time, you have \[\frac{dA}{dt}=2\pi r\frac{dr}{dt}\] You're given that \(\dfrac{dA}{dt}=1800\) and you want to find \(\dfrac{dr}{dt}\) when \(r=5000\).

OpenStudy (anonymous):

i keep on getting .115 km and im not exactly sure its corret

OpenStudy (anonymous):

You should be getting \[1800=2\pi(5000)\frac{dr}{dt}~~\iff~~\frac{dr}{dt}=\frac{1800}{10000\pi}=\frac{9}{50\pi}\] What does that get you?

OpenStudy (anonymous):

crap, i set it up incorrectly lol. thanks for the help :)

OpenStudy (anonymous):

You're welcome!